# CogSteer: Cognition-Inspired Selective Layer Intervention for Efficient Semantic Steering in Large Language Models

### v1

@@ Abstract
Despite their impressive capabilities, large language models (LLMs) often lack interpretability and can generate toxic content. While using LLMs as foundation models and applying semantic steering methods are widely practiced, we believe that efficient methods should be based on a thorough understanding of LLM behavior. To this end, we propose using eye movement measures to interpret LLM behavior across layers. We find that LLMs exhibit patterns similar to human gaze across layers and different layers function differently. Inspired by these findings, we introduce a heuristic steering layer selection and apply it to layer intervention methods via fine-tuning and inference. Using language toxification and detoxification as test beds, we demonstrate that our proposed CogSteer methods achieve better results in terms of toxicity scores while efficiently saving 97% of the computational resources and 60% of the training time. Our model-agnostic approach can be adopted into various LLMs, contributing to their interpretability and promoting trustworthiness for safe deployment.
CogSteer: Cognition-Inspired Selective Layer Intervention for Efficient Semantic Steering in Large Language Models
@@ 1 Introduction
Large language models (LLMs) Touvron et al. (2023); Achiam et al. (2023) have demonstrated strong abilities in natural language understanding and reasoning Zhao et al. (2023). Despite the continuous increase in performance and the emergence of new capabilities from scaling LLMs Wei et al. (2022), there are significant concerns about interpretability, safety, and trustworthiness in LLMs, leading to issues such as hallucinations Huang et al. (2023) and toxicity Schick et al. (2021). As the trend of using pre-trained LLMs as foundation models, followed by fine-tuning to align with human preferences Ouyang et al. (2022); Rafailov et al. (2024), the lack of transparency in LLMs has further limited the development of efficient post-training and inference methods.
@@
Figure 1: Framework of CogSteer. For an N-layer LLM, we first heuristically find the optimal layer M for semantic intervention. The upper block represents an adapter that is fine-tuned and inserted into the frozen layer M. The bottom block illustrates the operation of the attention module in layer M to steer the semantic direction towards safer outputs during inference.
@@
Previous work has introduced diverse methods of interpretability, including training linear classifiers as probes on top of hidden representations Belinkov (2022), projecting representations into vocabularies Geva et al. (2022), and intervening on the computation path, such as knowledge neurons Dai et al. (2022) and circuits Conmy et al. (2023); Ghandeharioun et al. (2024). However, these methods exhibit practical shortcomings. Probings rely on limited pre-defined classes, projection methods depend on human effort to interpret limited concepts, and interventions require well-designed prompts, making them difficult to scale and generalize. Furthermore, Bills et al. (2023) uses large language models to interpret smaller language models through natural language explanations, which we find problematic, as it involves using one black box to understand another.
We believe that a more promising approach to understanding LLMs is to analyze how they operate and behave similarly to the way humans process tokens and integrate information during reading. This can be achieved by using highly interpretable indicators from eye movement studies, such as fixation and regression Rayner (1998).
In this work, we introduce a new interpretability analysis method that utilizes eye movement data Luke and Christianson (2018); Hollenstein et al. (2020); Colman et al. (2022) collected by cognition researchers to analyze human reading behaviors. We correlate the hidden states of the FFNs in each layer with distinct eye movement measures and then observe the correlation values and their changes across layers. We reveal that LLMs show patterns similar to human gaze across layers, and that different layers function differently. In particular, the middle layers, which are responsible for further token processing and information integration, are promising venues to steer semantics in LLMs.
Based on the insights gained from our interpretability analysis using eye movement measures, we propose a heuristic steering layer selection approach to identify the optimal layer for semantic intervention, which can be widely applied during fine-tuning and inference. Specifically, to assess the efficiency of our method for semantic steering, we enhance the adapter tuning method by intervening at the optimal layer instead of using the last layer or all layers by default. Additionally, we propose an implicit layer contrastive intervention method during inference, which efficiently identifies and steers semantics toward safer generation directions.
Language toxification Dementieva et al. (2024) and detoxification Leong et al. (2023) are crucial semantic steering tasks, where toxification examines the jailbreak tendencies of LLMs, and detoxification aims for safer deployment of LLMs. Experiments on these tasks demonstrate that our proposed CogSteer method, for both fine-tuning and inference, is efficient—saving 97% of computational resources and 60% of training time—and achieves better performance compared with full-layer intervention using GPT-2 (+0.9%) and LLaMa2-7B (+2.8%) models. Moreover, our proposed implicit layer contrastive intervention method not only reduces the toxicity score from 87.2% to 59.1%, but also achieves a lower score compared to the original LLaMa2-7B model at 61.0%. Lastly, these consistent results echo the findings of our interpretability analysis.
Our contributions are summarized as follows: (1) We propose a novel interpretability method that utilizes eye movement measures to perform an in-depth analysis of LLM behavior across layers. (2) Building on these insights, we introduce a heuristic steering layer selection method and apply it efficiently to semantic steering in fine-tuning and inference. (3) Through extensive experimentation and analysis, we validate the effectiveness of our proposed method on language toxification and detoxification tasks, contributing to the deployment of explainable and trustworthy foundation models.
@@ 2 Related Work
Interpretability of Large Language Models
Interpretability research is essential for uncovering the mechanisms of LLMs, ensuring their safe and trustworthy deployment as foundation models Zhao et al. (2024). Various studies have analyzed how knowledge is stored in LLMs, focusing on concepts like knowledge neurons Dai et al. (2022) and the circuits Conmy et al. (2023); Yao et al. (2024). Knowledge neurons serve as a tool for localizing specific activated neurons, while the circuits framework identifies information propagation paths Goldowsky-Dill et al. (2023); Stolfo et al. (2023). We argue that these methods are insufficient as they rely on projections into the vocabulary space Rai et al. (2024) and well-designed prompts Bills et al. (2023), using humans to interpret limited samples as concepts Geva et al. (2022). Moreover, Wang et al. (2024a) preliminarily reveals that GPT predicts tokens more similarly to humans compared with shallow language models but lacks engagement with LLMs and a broad discussion on internal mechanisms. Unlike previous works, our method explores the interpretability of LLMs through human-interpretable indicators involving eye movement theory, to precisely understand and control model behavior (see § 3.13.2).
@@
Parameter efficient fine-tuning (PEFT) Han et al. (2024); He et al. (2022); Wang et al. (2023) has gained popularity due to the large number of frozen parameters in LLMs for generality, while adding only a few trainable parameters per task. Adapter tuning methods Houlsby et al. (2019); Poth et al. (2023) introduce new modules between layers of LLMs, whereas prompt tuning methods Liu et al. (2022) train prompt embeddings instead. Additionally, LoRA Hu et al. (2022) improves efficiency and helps mitigate overfitting. Our method for steering semantics via fine-tuning enhances adapter tuning by incorporating a selective layer intervention strategy, saving computational resources while delivering superior performance (see § 3.3). Moreover, steering semantics during inference presents an even more efficient approach. Contrastive decoding Li et al. (2023); Sennrich et al. (2024) guides the generation direction by comparing two output distributions. Wang et al. (2024b) proposed using instruction contrast to mitigate hallucinations. In contrast, our proposed implicit layer contrastive intervention efficiently identifies and steers the semantics of LLMs toward safe directions during inference (see § 3.4).

### v2

# CogSteer: Cognition-Inspired Selective Layer Intervention for Efficiently Steering Large Language Models

@@ Abstract
Large Language Models (LLMs) achieve remarkable performance through pretraining on extensive data. This enables efficient adaptation to diverse downstream tasks. However, the lack of interpretability in their underlying mechanisms limits the ability to effectively steer LLMs for specific applications. In this work, we investigate the intrinsic mechanisms of LLMs from a cognitive perspective using eye movement measures. Specifically, we analyze the layer-wise correlation between human cognitive indicators and LLM representations. Building on these insights, we propose a heuristic approach for selecting the optimal steering layer to modulate LLM semantics. To this end, we introduce an efficient selective layer intervention based on prominent parameter-efficient fine-tuning methods, which conventionally adjust either all layers or only the final layer. Additionally, we present an implicit layer contrastive intervention during inference to steer LLMs away from toxic outputs. Extensive experiments on natural language understanding, reasoning, and generation tasks, conducted on GPT-2, LLaMa2-7B, and Mixtral-7B, demonstrate the effectiveness and efficiency of our approach. As a model-agnostic framework, it enhances the interpretability of LLMs while improving efficiency for safe deployment.
CogSteer: Cognition-Inspired Selective Layer Intervention for Efficiently Steering Large Language Models
@@ 1 Introduction
Figure 1: Demonstration of CogSteer Intervention. For an N-layer LLM, we first heuristically find the optimal layer M for semantic intervention. The upper block represents an adapter that is fine-tuned and inserted into the frozen layer M. The bottom block illustrates the operation of the attention module in M to steer the semantic direction towards safer outputs during inference.
Large Language Models (LLMs) Dubey et al. (2024); Yang et al. (2024); Guo et al. (2025) have demonstrated strong capabilities in natural language understanding and reasoning Wei et al. (2022); Zhao et al. (2023) through pretraining on large datasets, followed by instruction tuning and alignment with human values Wei et al. (2021); Ouyang et al. (2022). Consequently, LLMs achieve excellent performance on downstream tasks with fine-tuning. However, their lack of interpretability and transparency limits the development of efficient fine-tuning and inference methods.
@@
To understand intrinsic mechanisms of LLMs, previous work has introduced various interpretability methods, including training linear classifiers as probes on top of hidden representations Belinkov (2022), projecting representations into vocabularies Geva et al. (2022), and intervening in the computation path, such as knowledge neurons Dai et al. (2022) and circuits Conmy et al. (2023); Ghandeharioun et al. (2024). However, these methods, which focus on a limited set of predefined classes, concepts, or prompts, have practical limitations in terms of scalability and generalization.
In this work, we introduce a novel interpretability analysis method that leverages eye movement data Luke and Christianson (2018); Hollenstein et al. (2020); Colman et al. (2022) collected by cognitive researchers to study human reading behaviors. Through correlation experiments, we find that LLM hidden states exhibit a strong correlation with human gaze, peaking in the middle layers. Using eye movement measures such as fixation and regression Rayner (1998) as human-interpretable indicators, we observe a hierarchical progression in LLMs, from initial syntactic and semantic processing to deeper integration and final prediction. Additionally, a comparison of correlation results between natural reading and task-specific reading suggests that the upper layers of LLMs are more capable of reasoning. Furthermore, advanced LLMs such as LLaMa, which incorporate instruction tuning and reinforcement learning from human feedback (RLHF), demonstrate enhanced reasoning capabilities, even in the middle layers.
LLM intervention Poth et al. (2023); Dong et al. (2024) refers to fine-tuning or applying inference methods to steer their semantics to align with specific tasks and data distributions. In addition to enhancing the understanding of LLM behavior, our interpretability analysis reveals that different layers serve distinct functions, with middle layers playing a crucial role in deeper syntactic and semantic processing. This enables us to first identify the most suitable layer for intervention, thereby improving task-specific performance. Based on these insights, we propose a heuristic approach for selecting the optimal steering layer for semantic intervention.
To achieve this, we refine prominent PEFT methods, which traditionally adjust either all layers or only the last layer. As shown in Figure 1, our proposed CogSteer framework first identifies the most suitable layer $M$ for semantic intervention based on the task. Instead of fine-tuning all layers or only the last layer, our method fine-tunes only the selected layer, enabling LLMs to better adapt to specific tasks and datasets. The number of learnable parameters in CogSteer is significantly reduced, requiring only $1/N$ of the parameters in LLMs, thereby improving parameter efficiency. Furthermore, we propose an implicit layer contrastive intervention method during inference, which efficiently identifies and steers semantics toward safer generation directions to evaluate the effectiveness of our proposed selective layer intervention.
Through extensive evaluations across diverse tasks and datasets, we demonstrate that the proposed selective layer intervention method achieves comparable or even superior performance with fewer parameters compared to the full-layer intervention baseline. Specifically, we observed an average absolute improvement of +1.7 on the GLUE benchmark for LLaMa2-7B, and an average absolute improvement of +5.8 on the GLUE benchmark for Mistral-7B, with only 3.1% of the parameters involved in full-layer fine-tuning. Moreover, in experiments on generation tasks, language toxification Dementieva et al. (2025), and detoxification Leong et al. (2023), our method achieves a +1.85% improvement in toxification compared to full-layer intervention and a +13.45% improvement in detoxification as compared to last-layer intervention.
Our main contributions are as follows:
(1) We are the first to propose leveraging eye movement measures to analyze LLM behavior. We publicly release the probing code to facilitate further research on interpretability from a cognitive perspective. (2) Through correlation analysis, we demonstrate a hierarchical progression in LLMs and introduce a heuristic steering layer selection method for efficient layer intervention. (3) Extensive experiments validate the effectiveness of our proposed method across various language understanding, reasoning, and generation tasks, contributing to the development of efficient and explainable foundation models.
@@ 2 Related Work
Interpretability
research is essential for uncovering the mechanisms of LLMs and ensuring their safe and trustworthy deployment as foundation models. Various studies have analyzed how knowledge is stored in LLMs Goldowsky-Dill et al. (2023); Stolfo et al. (2023); Rai et al. (2024); Bills et al. (2023); Geva et al. (2022), focusing on concepts such as knowledge neurons Dai et al. (2022) and circuits Conmy et al. (2023); Yao et al. (2024). Moreover, Wang et al. (2024a) reveal that GPT-2 predicts tokens more similarly to humans than shallow language models but lacks engagement with LLMs. Unlike previous works, our method investigates the interpretability of LLMs through human-interpretable indicators based on eye movement theory, enabling a more precise understanding and control of model behavior (see § 3 and  4.1).
@@
PEFT Han et al. (2024); Wang et al. (2023), including Adapter Houlsby et al. (2019); Poth et al. (2023) and LoRA Hu et al. (2022), has gained popularity due to its ability to maintain a large number of frozen parameters in LLMs for generality while introducing only a small number of trainable parameters per task. Our method for semantic steering via fine-tuning enhances PEFT methods by incorporating a selective layer intervention strategy, reducing computational costs while achieving superior performance (see § 4.2). Furthermore, steering semantics during inference offers an even more efficient approach. Contrastive decoding Li et al. (2023); Sennrich et al. (2024); Wang et al. (2024b) guides the generation process by comparing two output distributions. In contrast, our proposed implicit layer contrastive intervention efficiently identifies and steers the semantics of LLMs toward safe directions during inference (see § 4.3).
3 From Human Gaze to LLM Behavior
Recent studies Geva et al. (2021); Schuster et al. (2022) suggest that feed-forward networks (FFNs) function similarly to neural memory networks, capturing syntactic and semantic features as well as factual knowledge Chuang et al. (2023); Zhang et al. (2024). Meanwhile, research in cognitive science Rayner (1998) has shown that eye movement measures provide insights into the time required for human readers to process syntax, semantics, and integrate information. Motivated by these findings, we leverage eye movement measures to analyze their correlation with the hidden states of FFNs across different layers of LLMs.
Models. Correlation studies are conducted on the GPT-2 model Radford et al. (2019) at different sizes (12-layer small, 24-layer medium, 36-layer large) and the 32-layer LLaMa2-7B Touvron et al. (2023). Both GPT-2 and LLaMa2-7B use a decoder-only transformer architecture. Comparing earlier LLMs, such as GPT-2, with more advanced models like LLaMa2 provides valuable insights into their similarities and differences while also enhancing the robustness and generalizability of our findings.
Eye-movement Experiments and Datasets. We conduct a correlation analysis under two experimental conditions: natural reading and task-specific reading. For natural reading, we use the Provo Luke and Christianson (2018), GECO Colman et al. (2022), and ZuCo 2.0 Hollenstein et al. (2020) datasets. The ZuCo 2.0 dataset also includes task-specific reading experiments. In task-specific reading, participants are required to determine whether a specific relation type is present in a sentence. Relation detection is a high-level semantic and reasoning task that involves complex cognitive processing.
Correlation Analysis. Let $S_{j}$ denote the $j$-th sentence, consisting of $n_{j}$ words $w_{1},w_{2},\dots,w_{n_{j}}$. For each word $w_{i}$ in sentence $S_{j}$, we consider five eye movement measures: $e_{i}^{(k)},k\in\{\mathrm{\textit{std}},\,\mathrm{\textit{ffd}},\,\mathrm{\textit{gd}},\,\mathrm{\textit{trt}},\,\mathrm{\textit{gpt}}\}$, where each measure represents a scalar value. The hidden state at layer $l$ of the LLM for word $w_{i}$ is denoted as $\mathbf{h}_{l,i}\in\mathbb{R}^{d}$, where $d$ is the dimensionality of the hidden states.
To analyze the relationship between eye movement measures and LLM hidden states, we compute the Pearson correlation between each eye movement measure and the corresponding hidden states at each layer. Specifically, we concatenate hidden states and eye movement measures across all words in the dataset and apply principal component analysis to obtain a scalar representation of hidden states, ensuring alignment with eye movement measures (see Appendix D for details). The correlation is then defined as:
$\rho_{l,k}=\frac{\sum\limits_{i=1}^{n_{\textit{total}}}\bigl{(}h_{l,i}-\bar{h}_{l}\bigr{)}\bigl{(}e_{i}^{(k)}-\bar{e}^{(k)}\bigr{)}}{\sqrt{\sum\limits_{i=1}^{n_{\textit{total}}}\bigl{(}h_{l,i}-\bar{h}_{l}\bigr{)}^{2}}\sqrt{\sum\limits_{i=1}^{n_{\textit{total}}}\bigl{(}e_{i}^{(k)}-\bar{e}^{(k)}\bigr{)}^{2}}},$ (1)
where $h_{l,i}$ represents the processed hidden states aligned with the eye movement measures, and $\bar{h}_{l}$ and $\bar{e}^{(k)}$ are their respective means.
Results and Finding 1. Based on the correlation calculation in Eq. 1, we first analyze the correlation results for the natural reading task. To facilitate interpretation, we divide the layers of both models into three equal groups: premature, middle, and mature. Figures 2 (upper) and  7 present the correlation results between various eye movement measures and the hidden states of the LLMs, illustrating how these values evolve across layers.
The results indicate that the hidden states of different LLM layers exhibit a clear and strong correlation with human gaze, peaking in the middle bucket and reaching a secondary peak in the mature bucket. This trend is consistent across different eye movement measures and LLMs with varying layer sizes. Considering the nature of these eye movement measures, the increase in correlation in the premature bucket suggests that LLMs begin processing tokens by integrating syntactic and semantic features, reflecting an initial focus on token processing. In the middle bucket, the further increase in correlation signifies deeper syntactic and semantic processing, with the peak indicating the integration of linguistic features. In the mature bucket, the secondary peak likely reflects the final integration of information for word prediction.
The overall trend across the three buckets is similar for GPT-2 and LLaMa2-7B. For a detailed analysis of eye movement measures and additional findings, please refer to Appendix D.
Finding 1. (Layer-wise functionality) LLM hidden states exhibit a strong correlation with human gaze, peaking in the middle bucket and again in the mature bucket. This pattern suggests a hierarchical progression from initial syntactic and semantic processing to deeper integration and final prediction.
Results and Finding 2. Empirically, LLMs are trained with the next-token prediction objective. Modern LLMs demonstrate strong language understanding and reasoning abilities, raising the question: Are LLMs merely next-token predictors, or are they task reasoners? Figure 2 presents a comparison between natural reading and task-specific reading. The correlation patterns suggest that LLMs function as both next-token predictors and reasoners, as the trends in task-specific reading closely resemble those observed in natural reading.
Notably, for both GPT-2 and LLaMa2-7B models, the correlation values in the middle bucket during task-specific reading are higher than those observed during natural reading. In particular, for the LLaMa2-7B model, these values remain consistently higher in task-specific reading. We hypothesize that this indicates LLaMa2-7B is better suited for reasoning tasks and that the layers in the middle bucket and mature bucket are activated when processing complex tasks, as it is trained on a larger text corpus and incorporates more advanced post-training techniques.
Finding 2. (LLM functions as both next-token predictor and task reasoner) LLMs function as both next-token predictors and reasoners, with overall correlation trends aligning with human cognition indicators. Advanced training methods enhance their ability to reason and handle complex tasks.